OpenStudy (anonymous):

how do you obtain the graph of y=sqrt x+4-1from y=sqrtx

OpenStudy (anonymous):

is that 4-1 in the question ?

OpenStudy (anonymous):

if yes then just move the graph up by 3 units

OpenStudy (anonymous):

\[How/ to/ obtain /the/ graph/ of/ y=\sqrt{x+4}-1 from y=\sqrt{x}\]

OpenStudy (mertsj):

Move it 4 left and 1 down

OpenStudy (anonymous):

The -1 isn't inside of the sqrt sign

OpenStudy (mertsj):

And your point is?

OpenStudy (anonymous):

doesn't that make a difference or no

OpenStudy (mertsj):

No. If it were under the sqrt sign, we would subtract it from the 4 and the function would be different.

OpenStudy (anonymous):

oooo thanks a bunch

OpenStudy (mertsj):

yw